OpenStudy (anonymous):

what's the equation of the ellipse that has its center at the origin, the major axis is twice the length of the minor axis and (1, sqrt2) lies on the ellipse?

OpenStudy (anonymous):

i don't know how to incorporate the major/minor axis fact...

OpenStudy (anonymous):

minor axis=b major =2b x^2/(b)^2+(y^2)/(2b)^2=1 putting (1,sqrt(2) 1/b^2+2/(4b^2)=1 3/(2b^2)=1 b^2=3/2 b=sqrt(3/2) now u can find the equation putting sqrt(3/2) instead of b

OpenStudy (anonymous):

oh my gosh! it was that easy... thanks for clarifying!